You are thinking about investing in stock in a company that paid a dividend of $10 this year and whose dividends you expect will grow at 4 percent a year. The risk-free rate is 3 percent and you require a risk premium of 5 percent. If the price of the stock in the market is $200 a share, should you buy it?


1
Expert's answer
2021-09-06T18:54:57-0400

Yes

Making decision on buying share, it is important to calculate Share's fair value. Dividend discount model may be utilized in estimating the present stock value considering future cash flow.

"P =\frac{ D (1 + g)}{ r - g}"

D is the dividend, g is dividend's growth rate and R the needed return.

In this case, calculate r through adding risk premium of the stock and risk free rate. Hence;


"r = 3 + 5 = 8%" %

Present fair price of the stock,

P = 10 (1 + 4%) / 8% - 4% ⇒ $260 / share


This case is about comparing market price with fair price per share. The market price is 200 which is less than fair price, 260. Therefore, the share tends to be undervalued.